# manvr3d: A Platform for Human-in-the-loop Cell Tracking in Virtual Reality

### v2

@@ Abstract
We propose manvr3d, a novel VR-ready platform for interactive human-in-the-loop cell tracking. We utilize VR controllers and eye-tracking hardware to facilitate rapid ground truth generation and proofreading for deep learning-based cell tracking models.
Life scientists reconstruct the developmental history of organisms on the cellular level by analyzing 3D time-lapse microscopy images acquired at high spatio-temporal resolution. The reconstruction of such cell lineage trees traditionally involves tracking individual cells through all recorded time points, manually annotating their positions, and then linking them over time to create complete trajectories. Deep learning-based algorithms accelerate this process, yet depend heavily on manually-annotated high-quality ground truth data and curation. Visual representation of the image data in this process still relies primarily on 2D renderings, which greatly limits spatial understanding and navigation.
In this work, we bridge the gap between deep learning-based cell tracking software and 3D/VR visualization to create a human-in-the-loop cell tracking system. We lift the incremental annotation, training and proofreading loop of the deep learning model into the 3rd dimension and apply natural user interfaces like hand gestures and eye tracking to accelerate the cell tracking workflow for life scientists.
@@ 1 Introduction
Modern microscopes enable biologists to capture large scale 3D time-lapse datasets of embryonic development and other multi-cellular structures. Tracking of the imaged cells over time is a vital – yet non-trivial – task in the workflow of a life scientist studying the function and development of cells, tissues, and organisms. The result of the tracking process is a cell lineage tree (see manvr3d: A Platform for Human-in-the-loop Cell Tracking in Virtual Reality, right) that encodes information about the cellular ancestry.
@@
In this work, we present manvr3d (Multimodal ANnotations in Virtual Reality 3D, spoken as Manfred), extending the approach taken by ELEPHANT by bringing the annotation and linking steps into virtual reality (VR) and enabling users to perform these tasks with VR controllers and eye tracking. The central contributions of this work are:
@@
    Two implementations of VR-based cell tracking - One using handheld controllers in an interactive VR environment, and another using eye-tracking hardware in VR to accelerate the tracking process even further.
@@ 2 Related work
Various software solutions exist for analyzing biological data in VR. While ConfocalVR [20] or syGlass [15] offer general quantification and measurement tools, other tools like Arivis Pro VR111zeiss.com/microscopy/en/products/software/arivis-pro.html also offer features for cell tracking in volume data in VR. VR has also been used to segment biological 3D time-series data [5], to visualize the meshes resulting from this segmentation process [23] and to annotate those meshes [18]. Kaltenecker et al. [10] rely on Arivis Pro VR and syGlass to annotate cell segmentations, which they then use to train their 3D deep learning model DELiVR for automated cell segmentation to speed up the annotation process. Elor et al. developed the mixed reality environment BioLumin [3] to study the efficacy of crowd-sourced tissue annotation tasks, where the collected data is later used for training deep learning models.
In contrast to the aforementioned solutions, we present an open-source platform that enables researchers to integrate various kinds of natural user interfaces for human-in-the-loop cell tracking, resulting in a more streamlined and ergonomic process for both the creation of ground truth data, and the final proofreading step.
@@ 3.1 Software ecosystem
  • Sciview [6] and its underlying rendering framework scenery [7], to allow visualisation of large volumetric data together with 3D meshes. Both tools support a variety of natural user interfaces, such as VR headsets, or eye tracking hardware.
@@ 3.3 Organization and Data structure
manvr3d orchestrates all connections between the different components and ensures data consistency between the 2D data representation in Mastodon and the 3D representation in sciview. The data flow is visualized in Fig. 1 with color coded components. manvr3d (green) handles track editing events bidirectionally from either side and updates the other side accordingly. Time point changes are also communicated across components to maintain synchronized 2D (blue) and 3D (orange) views. Detailed information about Mastodons internal data structure are provided in Section A.1. We offer a graphical user interface that allows adjustment of various visualization parameters as well as launching a VR session (purple), either with or without additional hardware, such as eye trackers. The functionality of ELEPHANT (turquoise) is integrated as a menu inside the VR environment, enabling access to the training and prediction steps for the incremental deep learning model.
@@
Figure 2: Cell trajectories of the early development of a Tribolium castaneum (flour beetle), data provided by the Cell Tracking Challenge [12].
@@ 4 Cell Tracking with Natural User Interfaces
We provide two implementations for tracking cells in VR with natural user interfaces: using controllers (and in the future possibly hand gestures) and an experimental implementation for utilizing eye tracking hardware, where the user’s gaze directions are analyzed and cell tracks are created from this information. For intuitive interaction with the VR environment, we drew inspiration from popular VR productivity and creativity software like Gravity Sketch222gravitysketch.com or Shapelab333shapelabvr.com. Common one- and two-handed gestures for moving the observer, as well as scaling, rotating and translating of the dataset are implemented. A detailed controller layout can be found in Section A.2. Clicking a spot highlights it, and allows the user to either move it to a new position or delete it. In the same manner, an arbitrary number of new spots can be placed to annotate all cells in the current time point.
@@ 4.1 Cell Tracking with Handheld Controllers
A 3D cursor in form of a small sphere is attached to the right VR controller and allows interaction with the VR environment. Cell position annotation is performed by moving the 3D cursor into the target cell and pressing the right trigger button. Through the semi-transparent rendering of the image data, it is possible to precisely position the cursor inside the desired cell. With each annotation click, time is automatically advanced, making the tracing of a cell very rapid, by repeatedly clicking into the cell. By default, time advances backwards, as this is found to simplify handling of cell division events. This process is repeated until either the first time point is reached or the user manually terminates the tracking process by pressing the right B button. While tracking is active, a green placeholder track is shown to the user for immediate visual feedback, as can be seen in Fig. 3.
Figure 3: VR user tracking a cell with the controller. The active track is highlighted in green. The blue 3D cursor is partially obstructed by the volume. Data courtesy of William Lemon, HHMI Janelia Research Campus [11].
@@
It is possible to merge an active track into an existing spot – a cell division, since the animation is played backwards – by simply clicking on the target spot. The opposite is also possible: Clicking an existing spot with the trigger button to start a new track will use that spot as its origin point. Using both of these features in conjunction thus allows for bridging holes in existing tracks.
@@ 4.2 Cell Tracking via Eye Tracking
Tracking cells using gaze interaction has been explored in Bionic Tracking [4] by Günther et al. We incorporate this technique into our platform. While following a moving cell through the 3D volume data with their gaze, we record the user’s gaze directions and sample the volume at a uniform interval along each gaze ray. Finding the position of a cell along the spine thus turns into a 1-dimensional problem. The track is then reconstructed with a variant of the A* algorithm [22] that connects the closest local maxima from each gaze ray.
To avoid the Midas touch problem [8], it is important to remove any visual distractions that could lead the user to unintentionally look away from their target cell. For this reason, we implemented a dual input approach for cell tracking by eye tracking: as soon as the user is ready to start tracking a cell with their eyes, they press the left trigger button. This starts playback of the dataset and the continuous collection of gaze directions and volume density samples along each ray. Once the user interrupts the tracking with the trigger button, or if the first time point is reached, the collected gaze rays and their sampled values are analyzed and subsequently sent to Mastodon. A collection of rays is plotted in Fig. 4, where each ray originates from the positive Y axis and extends downwards. Changes over time are plotted along the X axis. We apply a subtle Gaussian smoothing to extract local maxima more robustly.
Figure 4: A collection of gaze rays across time, collected by following a moving cell with one’s eyes. We slide a simple Gauss kernel in the shape of [0.25,0.5,0.25]0.250.50.25[0.25,0.5,0.25][ 0.25 , 0.5 , 0.25 ] along each ray to smooth the signal and extract local maxima more effectively. Here we show the effect of various iterations of Gauss smoothing. It can be seen that no or low amounts of smoothing can lead to incorrectly extracted tracks, and starting from 4 iterations onward we are able to extract the correct cell track.
5 Performance
@@
Both primitives types of the 3D representation, spheres for spots and cylinders for links, are currently rendered as instanced meshes on the GPU. Two instance pools are populated during the initialization phase – one for spots and one for links. We found that instance pre-generation is faster than on-the-fly generation. The instances in these pools are then positioned and colored according to the current time point. If a time point requires more spots, they are dynamically allocated to the pool.
We benchmarked sciview’s capability to render geometry instances (see Table 1) with two differently-sized datasets and found that for up to several ten thousand cells, there is no negative effect on performance [^4] The frame rates refer to a render resolution of 1920x1080 pixels, on an XMG Fusion 15 laptop with NVIDIA RTX 4070 GPU and Intel Core i9-14900HX processor, running Windows 10 build 19044.5371, JDK version 21.0.5, NVIDIA driver version 566.36.. Turning off the overlayed volume rendering yields slightly faster frame rates.
Table 1: Frame rate and scene population time depending on the number of geometry instances rendered in the scene. The scene population time is only required once during scene initialization.
@@
Frame rate with volume (Hz) 185 29
Frame rate without volume (Hz) 230 32
@@ 6 Summary and future work
In this work, we presented manvr3d, a platform to allow VR/natural user interface-based cell tracking together with two example implementations, using handheld controllers or eye tracking hardware. We couple this annotation process with the ELEPHANT deep learning model for rapid training data acquisition and provide track editing tools for proofreading of the predictions. manvr3d easily scales to several ten thousand cells.
@@
To quantify the potential improvement of our human-in-the-loop tracking approach over conventional methods, we plan to conduct a user study that compares the speed and accuracy of fully manual methods with automated methods and the approach taken in this project.
@@ Acknowledgements.
The authors thank Ko Sugawara for his ongoing support with the ELEPHANT integration. Thanks to Vladimír Ulman and Ruoshan Lan for their work on a prototype of this project. The authors also want to thank Elias Barriga, Valentin Ruffine, Ana Patricia Ramos and Jaime Hidalgo for their feedback on the software. This work was partially funded by the Center for Advanced Systems Understanding (CASUS), financed by Germany’s Federal Ministry of Education and Research (BMBF) and by the Saxon Ministry for Science, Culture and Tourism (SMWK) with tax funds on the basis of the budget approved by the Saxon State Parliament. This work was also funded by the French National Research Agency (France BioImaging, ANR-24-INBS-0005 FBI BIOGEN).
Appendix A Appendix
@@ A.1 Mastodon data structure
Spot colors can be customized in Mastodon using tag sets; this aids visual analysis of the dataset. The bridge supports transferring these colors to the 3D tracks.
@@ A.2 VR interactions
Figure 5: VR controller layout for a pair of Meta Quest 2 controllers.

### v3

@@ Abstract
We propose manvr3d, a VR platform for immersive, AI-assisted human-in-the-loop cell tracking. Life scientists reconstruct the developmental history of organisms at the cellular level by analyzing 3D time-lapse microscopy images acquired at high spatio-temporal resolution. However, reconstruction of cell trajectories and lineage trees is a highly time consuming and error prone task. Common tools are often limited to 2D image display, which greatly limits spatial understanding and navigation. Deep Learning-based algorithms accelerate this process, yet depend heavily on manually-annotated, high-quality ground truth data and curation. In this work, we bridge the gap between Deep Learning-based cell tracking software and 3D/VR visualization to create a hybrid AI-human-in-the-loop cell tracking system. We lift the incremental annotation, training and proofreading loop of the deep learning model into the third dimension and apply natural user interfaces like hand gestures and eye tracking to accelerate the cell tracking workflow for life scientists. We present here the technical architecture of our platform and first analysis of performance. Our code is released open source.
@@ 1 Introduction
Modern microscopes enable biologists to capture large scale 3D time-lapse datasets of embryonic development and other multi-cellular structures. Tracking of the imaged cells over time is a vital—yet non-trivial—task in the workflow of a life scientist studying the function and development of cells, tissues, and organisms. The result of the tracking process is a cell lineage tree (see manvr3d: A Platform for Human-in-the-loop Cell Tracking in Virtual Reality, right) that encodes information about the cellular ancestry.
@@
In this work, we present manvr3d (Multimodal ANnotations in Virtual Reality 3D, pronounced “manfred”), extending the approach taken by ELEPHANT by bringing the annotation and linking steps into virtual reality (VR) and enabling users to perform these tasks with VR controllers and eye tracking. The central contributions of this work are:
@@
    Two implementations of VR-based cell tracking—One using handheld controllers in an interactive VR environment, and another using eye-tracking hardware in VR to accelerate the tracking process even further.
In this paper we describe the development of a functioning prototype. A full user study is beyond the scope of this work, in Section 5 however, we show indicative numbers for both rendering and annotation performance.
@@ 2 Related work
Various software solutions exist for analyzing and visualizing biological data in 3D and/or VR environments. For cell lineage data, visualization and annotation solutions exist for both 2D [13, 22] and 3D [9, 23]. Neither of these support immersive rendering in VR. Leeuw et al. [3] address this gap by rendering cell trajectories in a CAVE [^1] Cave automatic virtual environment system with superimposed volume time-series. None of those solutions integrate the visualization aspects into a wider analysis platform and do not offer integration of a machine learning model.
Additionally, proprietary solutions like ConfocalVR [25] or syGlass [18] offer general quantification and measurement tools for biological image data, Arivis Pro VR222zeiss.com/microscopy/en/products/software/arivis-pro.html also offers features for cell tracking in volume data in VR. Arivis and syGlass have been utilized by Kaltenecker et al. [12] to annotate cell segmentations, which they then use to train their 3D deep learning model DELiVR for automated cell segmentation to speed up the annotation process. Note that in this work, no cell tracking is performed. Elor et al. developed the mixed reality environment BioLumin [4] to study the efficacy of crowd-sourced tissue annotation tasks, where the collected data is later used for training deep learning models. VR has been used in [5] to segment biological 3D time-series data, [28] visualized meshes resulting from the segmentation process, and [21] further annotate those meshes in VR. Again, no cell tracking is performed.
In contrast to the aforementioned solutions, we present an immersive open-source visualization platform for human-in-the-loop cell tracking with natural user interfaces to enable a streamlined and ergonomic process for both the creation of ground truth data and the final proofreading step. Our system is embedded into the Fiji [24] ecosystem, resulting in easier adoption and integration into existing pipelines.
@@ 3.1 Software ecosystem
  • Sciview [6] and its underlying rendering framework scenery [8], to allow visualization of large volumetric data together with 3D meshes. Both tools support a variety of natural user interfaces, such as VR headsets, or eye tracking hardware.
@@ 3.3 Organization and Data structure
manvr3d orchestrates all connections between the different components and ensures data consistency between the 2D data representation in Mastodon and the 3D representation in sciview. The data flow is visualized in Fig. 1 with color coded components. manvr3d (green) handles track editing events bidirectionally from either side and updates the other side accordingly. Time point changes are also communicated across components to maintain synchronized 2D (blue) and 3D (orange) views. Detailed information about Mastodons internal data structure are provided in Supplement A.1. We offer a graphical user interface that allows adjustment of various visualization parameters as well as launching a VR session (purple), either with or without additional hardware, such as eye trackers. The functionality of ELEPHANT (turquoise) is integrated as a menu inside the VR environment, enabling access to the training and prediction steps for the incremental deep learning model.
@@ 4 Cell Tracking with Natural User Interfaces
We provide two implementations for tracking cells in VR with natural user interfaces: using controllers (and in the future possibly hand gestures) and an experimental implementation for utilizing eye tracking hardware, where the user’s gaze directions are analyzed and cell tracks are created from this information. For intuitive interaction with the VR environment, we drew inspiration from popular VR productivity and creativity software like Gravity Sketch333gravitysketch.com or Shapelab444shapelabvr.com. Common one- and two-handed gestures for moving the observer, as well as scaling, rotating and translating of the dataset are implemented. A detailed controller layout can be found in Supplement A.2. Clicking a spot highlights it, and allows the user to either move it to a new position or delete it. In the same manner, an arbitrary number of new spots can be placed to annotate all cells in the current time point.
@@ 4.1 Cell Tracking with Handheld Controllers
A 3D cursor in form of a small sphere is attached to the right VR controller and allows interaction with the VR environment. Cell position annotation is performed by moving the 3D cursor into the target cell and pressing the right trigger button. Through the semi-transparent rendering of the image data, it is possible to precisely position the cursor inside the desired cell. With each annotation click, time is automatically advanced, making the tracing of a cell very rapid, by repeatedly clicking into the cell. By default, time advances backwards, as this is found to simplify handling of cell division events. This process is repeated until either the first time point is reached or the user manually terminates the tracking process by pressing the right B button.
Figure 2: VR user looking at a Platynereis dataset, annotated with controllers and ELEPHANT. Data courtesy of Tomancak Lab, MPI-CBG.
@@
It is possible to merge an active track into an existing spot—a cell division, since the animation is played backwards—by simply clicking on the target spot. The opposite is also possible: Clicking an existing spot with the trigger button to start a new track will use that spot as its origin point. Using both of these features in conjunction thus allows for bridging holes in existing tracks.
Figure 3: Left: Scheme of the graph search algorithm that connects the closest local maxima in subsequent gaze rays following the first local maximum found in the first ray. Right: A collection of gaze rays across time, collected by following a moving cell with one’s eyes. We slide a simple Gauss kernel in the shape of [0.25,0.5,0.25]0.250.50.25[0.25,0.5,0.25][ 0.25 , 0.5 , 0.25 ] along each ray to smooth the signal and extract local maxima more effectively. Here we show the effect of various iterations of Gauss smoothing. It can be seen that no or low amounts of smoothing can lead to incorrectly extracted tracks, and starting from 4 iterations onward we are able to extract the correct cell track.
@@ 4.2 Cell Tracking via Eye Tracking
Tracking cells using gaze interaction has been explored in Bionic Tracking [7] by Günther et al. We incorporate this technique into our platform. While following a moving cell through the 3D volume data with their gaze, we record the user’s gaze directions and sample the volume at a uniform interval along each gaze ray. Finding the position of a cell along this gaze ray thus turns into a 1-dimensional problem, because we can assume that the first local maximum along the ray corresponds to the target cell. After calculating the local maxima for each ray, the track is reconstructed with a variant of the A* algorithm [27] that connects the closest local maxima from subsequent gaze rays, see Fig. 3.
To avoid the Midas touch problem [10], it is important to remove any visual distractions that could lead the user to unintentionally look away from their target cell. For this reason, we implemented a dual input approach for cell tracking by eye tracking: as soon as the user is ready to start tracking a cell with their eyes, they press the left trigger button. This starts playback of the dataset and the continuous collection of gaze directions and volume density samples along each ray. Once the user interrupts the tracking with the trigger button, or if the first time point is reached, the collected gaze rays and their sampled values are analyzed and subsequently sent to Mastodon. A collection of rays is plotted in Fig. 3, where each ray originates from the positive Y axis and extends downwards. Changes over time are plotted along the X axis.
5 Performance and Results
5.1 Rendering Performance
@@
Both primitive types of the 3D representation, spheres for spots and cylinders for links, are currently rendered as instanced meshes on the GPU. Two instance pools are populated during the initialization phase—one for spots and one for links. We found that instance pre-generation is faster than on-the-fly generation. The instances in these pools are then positioned and colored according to the current time point. If a time point requires more spots, they are dynamically allocated to the pool.
We benchmarked manvr3d’s capability to render geometry instances (see Table 1) with two differently-sized datasets and found that for up to several ten thousand cells, there is no negative effect on performance [^5] The frame rates refer to a render resolution of 1920x1080 pixels, on an XMG Fusion 15 laptop with NVIDIA RTX 4070 GPU and Intel Core i9-14900HX processor, running Windows 10 build 19044.5371, JDK version 21.0.5, NVIDIA driver version 566.36.. Turning off the overlaid volume rendering yields slightly faster frame rates.
Table 1: Frame rate and scene population time, depending on the number of geometry instances rendered in the scene. The scene population time is only required once during scene initialization.
@@
Frame rate with volume (fps) 185 29
Frame rate without volume (fps) 230 32
@@
5.2 Annotation performance
As stated before, performing a full user study is beyond the scope of this work. We instead performed two partial annotations (Table 2) of a Platynereis dataset and a Drosophila dataset [14], to compare annotation speeds between the Mastodon approach (2D), VR-based controller tracking (VR) and ELEPHANT (+DL). We measured the time taken to create 10 tracks each. We then used the pre-trained versatile ELEPHANT model and trained it over 5 epochs on the annotations created, and divided the final amount of tracks (104 for Platynereis and 74 for Drosophila) by the sum of training and prediction time. The training/prediction time is independent of the manual annotation method used.
Table 2: Annotation times for 2D, VR and Deep Learning tracking.
Dataset Size Type Time/track
Platynereis 700x660x113, 101 time points 2D 3.85 min
VR 0.65 min
+DL 0.125 min
Drosophila 151x101x29, 31 time points 2D 1.07 min
VR 0.16 min
+DL 0.02 min
These numbers indicate that VR-based tracking can significantly outperform manual 2D annotations, by a factor of about 6. Although the time needed for prediction and training in the ELEPHANT model remains the same for both methods, manvr3d can provide a significant benefit for the annotation aspect.
@@ 6 Summary and future work
In this work, we presented manvr3d, a platform to allow VR/natural user interface-based cell tracking together with two example implementations, using handheld controllers or eye tracking hardware. We couple this annotation process with the ELEPHANT deep learning model for rapid training data acquisition and provide track editing tools for proofreading of the predictions. manvr3d easily scales to several thousand cells.
@@
To quantify the improvement—indicated in Section 5.2—of our human-in-the-loop tracking approach over conventional methods, we plan to conduct a user study that compares the speed and accuracy of fully manual methods with automated methods and the approach taken in this project, using a variety of datasets.
7 Software Availability
@@ Acknowledgements.
The authors thank Ko Sugawara for his ongoing support with the ELEPHANT integration. Thanks to Vladimír Ulman and Ruoshan Lan for their work on a prototype of this project, and to Jan Tiemann for providing code and support for 3D user interface elements. The authors also want to thank Elias Barriga, Valentin Ruffine, Ana Patricia Ramos and Jaime Hidalgo for their feedback on the software. SP was partially funded by the Center for Advanced Systems Understanding (CASUS), financed by Germany’s Federal Ministry for Research, Technology, and Space (BMFTR) and by the Saxon Ministry for Science, Culture and Tourism (SMWK) with tax funds on the basis of the budget approved by the Saxon State Parliament. JYT was funded by the French National Research Agency (France BioImaging, ANR-24-INBS-0005 FBI BIOGEN). UG was supported by GoBio Initial, 03LW0622, by the Germany’s Federal Ministry for Research, Technology, and Space (BMFTR).
Appendix A Supplementary Material
Figure 4: VR controller layout for a pair of Meta Quest 2 controllers.
@@ A.1 Mastodon data structure
Spot colors can be customized in Mastodon using tag sets; this aids visual analysis of the dataset. manvr3d supports transferring these colors to the 3D tracks.